# Importance-Weighted Non-IID Sampling for Flow Matching Models

### v1

@@ C.2 Evolution Along a Trajectory
As in Definition˜2, samples evolve according to
$\dot{X}_{t}^{(i)}=v\big(X_{t}^{(i)},t\big)+u\big(X^{(i)}_{t},X^{(-i)}_{t},t\big),\!\!\!\!\!\qquad X_{0}^{(i)}\overset{\text{IID}}{\sim}p_{0},$ (S28)
@@
$\frac{\mathrm{d}}{\mathrm{dt}}\log w_{\phi,t}(X_{t}^{(i)})$ $=\nabla_{x}\log w_{\phi,t}(X_{t}^{(i)})\cdot\dot{X}_{t}^{(i)}$ (S29)
@@
$\nabla_{x}\log w_{\phi,t}(x)$ $=\nabla_{x}\log p_{t}(x)-\nabla_{x}\log p^{\prime\prime}_{\phi,t}(x)$ (S30)
@@
Combining Equations˜S29, S27 and S30 yields
$\frac{\mathrm{d}}{\mathrm{dt}}\log w_{\phi,t}(X_{t}^{(i)})$ (S31)
@@
which matches to Theorem˜2.
@@ C.3 Evolution Along a Trajectory for Rectified Flows
For rectified flows, Appendix˜B yields
$s(x,t)=\frac{tv(x,t)-x}{1-t},$ (S32)
@@
$s^{\prime\prime}_{\phi}(x,t)=\frac{t\big(v(x,t)+r_{\phi}(x,t)\big)-x}{1-t}.$ (S33)
@@
$s^{\prime\prime}_{\phi}(x,t)-s(x,t)=\frac{t}{1-t}r_{\phi}(x,t).$ (S34)
Substituting Equations˜S32, S33 and S34 into Equation˜S31 gives
$\quad\frac{\mathrm{d}}{\mathrm{dt}}\log w_{\phi,t}(X_{t}^{(i)})$ (S35)
@@
which matches to Corollary˜1.
@@ C.4 Evolution at a Fixed Position for Rectified Flows
Combining Equations˜S25 and S26 with the rectified-flow identity in Appendix˜B, we obtain
$\quad\partial_{t}\log w_{t}(x)$ (S36)
@@ Appendix D Diversity Objectives Used in Experiments
  • Determinant Point Process (DPP) [kulesza2012determinantal]:
@@
  • Particle Guidance (PG) [DBLP:conf/iclr/CorsoXBBJ24]:
@@
where $K$ is defined in Equation˜11, $T_{r}$ is the $r$-th Chebyshev polynomial [mason2002chebyshev], and $K_{\mathrm{H}}$ is defined as
$K_{\mathrm{H}}=\tfrac{1}{\sum_{r=0}^{n-1}a_{r}}\sum_{r=0}^{n-1}a_{r}\,T_{r}(L_{\cos}),$ (S37)
with $a_{0}{=}1$, $a_{r}{=}2\!\left(1-\tfrac{r}{n}\right)$ when $r{\geq}1$, $L_{\cos,ij}{=}\langle\hat{x}^{(i)},\hat{x}^{(j)}\rangle$, and $\hat{x}^{(i)}{=}x^{(i)}/\|x^{(i)}\|$. Gradients are normalized in implementation, so constant scalings are omitted. For the strength parameter $\lambda$ of the diversity velocity (see Equation˜14), we set $\lambda=1$ for Sec.˜4.1, $\lambda=0.1$ for Sec.˜4.3, and $\lambda=0.3$ for Sec.˜4.4.
@@ Appendix E Experiment Implementation Details
For the Gaussian mixture experiment (Sec.˜4.1), we use 100 flow-matching steps with uniformly spaced time intervals and an Euler solver. For this task, we train the residual velocity for 100 epochs, sampling a random $t\in[0,1]$ for each example. We use the AdamW [DBLP:conf/iclr/LoshchilovH19] optimizer with a learning rate of $10^{-4}$, weight decay of $10^{-3}$, and a batch size of 1,0001{,}000.
For both text-to-image generation (Sec.˜4.3) and image inpainting (Sec.˜4.4), we use 28 uniformly spaced sampling steps with an Euler solver and disable Classifier-Free Guidance (CFG) [DBLP:journals/corr/abs-2207-12598] for all experiments. We compute diversity objectives using the spatial mean of latent features. For inpainting, we further adopt a fixed text prompt, “a realistic photo,” across all input images.
@@ Appendix F Construction of Ground-Truth Density
We obtain ground-truth densities for the non-IID marginal by repeatedly resampling from the non-IID dataset and pooling the draws into a single large set. Given a sample set $\mathcal{Y}=\{y_{i}\}_{i=1}^{B}$, $y_{i}\in\mathbb{R}^{d}$, we evaluate densities at each sample $x\in\mathcal{Y}$ using a leave-one-out strategy—excluding $x$ from its own neighborhood—so that n=B1n{=}B{-}1 samples contribute to the estimate. Letting $I_{d}$ denote the $d{\times}d$ identity, we follow prior work [Loader1996, HjortJones1996] and estimate $\log p^{\prime}(x)$ by fitting a local quadratic model to $\log p^{\prime}$ using kernel-weighted moments computed across multiple kNN-defined bandwidths (scales $h$). We then combine the resulting per-scale log-density estimates via inverse-variance weighting.
@@
$\mu=\frac{m_{1}}{m_{0}},\qquad S=\frac{M_{2}}{m_{0}}-\mu\mu^{\top}.$ (S42)
@@
$\widehat{p}^{\prime}_{h}(x)=\exp\{\widehat{a}_{h}(x)\},$ (S44)
@@
For the unit-bandwidth Gaussian kernel [WandJones1995], the kernel roughness is
$R(K)=\int K(u)^{2}\,\mathrm{d}u=(4\pi)^{-d/2}.$ (S45)
Using Kish’s effective sample size [Kish1965],
$n_{\mathrm{eff}}(x,h)=\frac{(\sum_{i}w_{i})^{2}}{\sum_{i}w_{i}^{2}},$ (S46)
a Wald approximation [wald1949note] leads to
$\mathrm{Var}(\widehat{a}_{h}(x))\approx\frac{R(K)}{\,n_{\mathrm{eff}}(x,h)\,h^{d}\,\widehat{p}^{\prime}_{h}(x)},\\$ (S47)
@@
$\widehat{a}(x)=\frac{\sum_{h}I_{h}(x)\,\widehat{a}_{h}(x)}{\sum_{h}I_{h}(x)},\qquad\widehat{p}^{\prime}(x)=\exp\{\widehat{a}(x)\}.$ (S49)
@@ Appendix G Baseline Density Estimators
kNN density [loftsgaarden1965nonparametric].
@@
$\widehat{p}_{\text{kNN}}(x)$ $=\;\frac{k}{B\,V_{d}\,\big(r_{k}(x)\big)^{d}}.$ (S50)
Gaussian KDE with Silverman’s rule [DBLP:books/sp/Silverman86].
@@
$\widehat{p}_{\text{KDE}}(x)$ $=\frac{1}{B}\sum_{j=1}^{B}\frac{\exp\!\Big(-\frac{\|x-x^{(j)}\|_{2}^{2}}{2h^{2}}\Big)}{(2\pi)^{d/2}h^{d}},$ (S51)
$h$ $=\sigma\left(\frac{4}{d+2}\right)^{\!\frac{1}{d+4}}\,B^{-\frac{1}{d+4}},$ (S52)
$\sigma$ $=\tfrac{1}{d}\sum_{\ell=1}^{d}\text{std}(X_{:\ell}).$ (S53)
Multivariate Gaussian fit (MGF) [DBLP:journals/jei/BishopN07].
@@
$\quad\log\widehat{p}_{\text{MGF}}(x^{(i)})$ (S55)
@@ Appendix H Metrics for Ranking
Spearman’s $\rho$ [spearman1987proof].
Spearman’s $\rho$ assesses the strength of a monotonic relationship between two rankings. Assign 1‑based ranks (1 = best) with average‑tied ranks and compute the Pearson correlation between predicted and ground‑truth ranks:
@@
Graded Average Precision (gAP) [DBLP:conf/sigir/MoffatM0A22].
Graded AP extends Average Precision to graded (multi‑level) relevance. First, sort the data by predicted scores. Normalize per query to $[0,1]$: shift if any grade is negative; then, if the maximum exceeds $1$, divide all grades by that maximum. Define
$\mathrm{Prec}@i$ $=\frac{\sum_{j=1}^{i}g_{j}}{i},\quad\mathrm{gAP}=\frac{\sum_{i=1}^{n}g_{i}\,\mathrm{Prec}@i}{\sum_{i=1}^{n}g_{i}}.$ (S57)
When $g_{i}\in\{0,1\}$, $\mathrm{gAP}$ reduces to standard AP. Larger values indicate that higher‑graded items are concentrated near the top of the ranking.
@@ Appendix I Illustrative Example of Joint ODE Trajectories
Figure S1 visualizes joint ODE trajectories on a 2D Gaussian mixture with three components of mixture weights $1/6$, $1/3$, and $1/2$. Under the IID baseline (no diversity objective), trajectories collapse to the highest-weight component, as expected. Introducing a DPP diversity objective yields Non-IID trajectories that collectively cover all three modes, improving coverage under a fixed trajectory budget. Augmenting DPP with our score-based diversity-velocity regularizer (SR) draws trajectories closer to the mode centers while preserving coverage (see the white rectangle). Finally, importance weighting (IW) assigns a weight to each trajectory. In the full method (DPP + SR + IW), the sample closest to the highest-weight component receives the largest weight (white circle) and that closest to the lowest-weight component receives the smallest, correcting the Non-IID sampling bias and yielding an unbiased estimator with respect to the IID target.
Figure S1: Joint ODE trajectories on a three-component Gaussian mixture. All methods start from identical initial states. “IID” is the baseline without a diversity objective and collapses to the highest-weight mode. “Non-IID (DPP)” uses a determinantal point process objective to encourage sample diversity, enabling the three trajectories to discover all three modes. “SR” denotes our score-based diversity-velocity regularization (the soft version is used in this example). Combining DPP + SR pulls samples toward the underlying modes while preserving coverage (white rectangle). “IW” denotes importance-weight estimation. The full method (DPP + SR + IW) further assigns an importance weight to each trajectory; larger markers indicate higher estimated weight (white circles). For visualization, arrow lengths are rescaled nonlinearly while preserving their relative ordering. The background shows the relative value of $\log p(x)$, where $p$ is the target density corresponding to the IID ODE; yellow indicates higher density and purple indicates lower density.

### v2

@@ C.2 Evolution Along a Trajectory
As in Definition 2, samples evolve according to
@@
$\frac{\mathrm{d}}{\mathrm{dt}}\log w_{\phi,t}(X_{t}^{(i)})$ $=\nabla_{x}\log w_{\phi,t}(X_{t}^{(i)})\cdot\dot{X}_{t}^{(i)}$ (55)
@@
$\nabla_{x}\log w_{\phi,t}(x)$ $=\nabla_{x}\log p_{t}(x)-\nabla_{x}\log p^{\prime\prime}_{\phi,t}(x)$ (56)
@@
Combining Eq (55), Eq (53), Eq (56) yields
$\frac{\mathrm{d}}{\mathrm{dt}}\log w_{\phi,t}(X_{t}^{(i)})$ (57)
@@
which matches to Theorem 2.
@@ C.3 Evolution Along a Trajectory for Rectified Flows
For rectified flows, Appendix B yields
$s(x,t)=\frac{tv(x,t)-x}{1-t},$ (58)
@@
$s^{\prime\prime}_{\phi}(x,t)=\frac{t\big(v(x,t)+r_{\phi}(x,t)\big)-x}{1-t}.$ (59)
@@
$s^{\prime\prime}_{\phi}(x,t)-s(x,t)=\frac{t}{1-t}r_{\phi}(x,t).$ (60)
Substituting Eq. (58), Eq. (59), Eq. (60) into 57 gives
$\quad\frac{\mathrm{d}}{\mathrm{dt}}\log w_{\phi,t}(X_{t}^{(i)})$ (61)
@@
which matches to Corollary (1).
@@ C.4 Evolution at a Fixed Position for Rectified Flows
Combining Eq. (51) and Eq. (52) with the rectified-flow identity in Appendix B, we obtain
$\quad\partial_{t}\log w_{t}(x)$ (62)
@@ Appendix D Diversity Objectives Used in Experiments
  • Determinant Point Process (DPP) [Kulesza et al., 2012]:
@@
  • Particle Guidance (PG) [Corso et al., 2024]:
@@
where $K$ is defined in Eq. (11), $T_{r}$ is the $r$-th Chebyshev polynomial [Mason and Handscomb, 2002], and $K_{\mathrm{H}}$ is defined as
$K_{\mathrm{H}}=\tfrac{1}{\sum_{r=0}^{n-1}a_{r}}\sum_{r=0}^{n-1}a_{r}\,T_{r}(L_{\cos}),$ (63)
with $a_{0}{=}1$, $a_{r}{=}2\!\left(1-\tfrac{r}{n}\right)$ when $r{\geq}1$, $L_{\cos,ij}{=}\langle\hat{x}^{(i)},\hat{x}^{(j)}\rangle$, and $\hat{x}^{(i)}{=}x^{(i)}/\|x^{(i)}\|$. Gradients are normalized in implementation, so constant scalings are omitted. For the strength parameter $\lambda$ of the diversity velocity (see Eq. (14)), we set $\lambda=1$ for Section (4.1), $\lambda=0.1$ for Section (4.3), and $\lambda=0.3$ for Section (4.4).
@@ Appendix E Experiment Implementation Details
For the Gaussian mixture experiment (Section 4.1), we use 100 flow-matching steps with uniformly spaced time intervals and an Euler solver. For this task, we train the residual velocity for 100 epochs, sampling a random $t\in[0,1]$ for each example. We use the AdamW [Loshchilov and Hutter, 2019] optimizer with a learning rate of $10^{-4}$, weight decay of $10^{-3}$, and a batch size of 1,0001{,}000.
For both text-to-image generation (Section 4.3) and image inpainting (Section 4.4), we use 28 uniformly spaced sampling steps with an Euler solver and disable Classifier-Free Guidance (CFG) [Ho and Salimans, 2022] for all experiments. We compute diversity objectives using the spatial mean of latent features. For inpainting, we further adopt a fixed text prompt, “a realistic photo,” across all input images.
@@ Appendix F Construction of Ground-Truth Density
We obtain ground-truth densities for the non-IID marginal by repeatedly resampling from the non-IID dataset and pooling the draws into a single large set. Given a sample set $\mathcal{Y}=\{y_{i}\}_{i=1}^{B}$, $y_{i}\in\mathbb{R}^{d}$, we evaluate densities at each sample $x\in\mathcal{Y}$ using a leave-one-out strategy—excluding $x$ from its own neighborhood—so that n=B1n{=}B{-}1 samples contribute to the estimate. Letting $I_{d}$ denote the $d{\times}d$ identity, we follow prior work [Loader, 1996, Hjort and Jones, 1996] and estimate $\log p^{\prime}(x)$ by fitting a local quadratic model to $\log p^{\prime}$ using kernel-weighted moments computed across multiple kNN-defined bandwidths (scales $h$). We then combine the resulting per-scale log-density estimates via inverse-variance weighting.
@@
$\mu=\frac{m_{1}}{m_{0}},\quad S=\frac{M_{2}}{m_{0}}-\mu\mu^{\top}.$ (68)
@@
$\widehat{p}^{\prime}_{h}(x)=\exp\{\widehat{a}_{h}(x)\},$ (70)
@@
For the unit-bandwidth Gaussian kernel [Wand and Jones, 1995], the kernel roughness is
$R(K)=\int K(u)^{2}\,\mathrm{d}u=(4\pi)^{-d/2}.$ (71)
Using Kish’s effective sample size [Kish, 1965],
$n_{\mathrm{eff}}(x,h)=\frac{(\sum_{i}w_{i})^{2}}{\sum_{i}w_{i}^{2}},$ (72)
a Wald approximation [Wald, 1949] leads to
$\mathrm{Var}(\widehat{a}_{h}(x))\approx\frac{R(K)}{\,n_{\mathrm{eff}}(x,h)\,h^{d}\,\widehat{p}^{\prime}_{h}(x)},\\$ (73)
@@
$\widehat{a}(x)=\frac{\sum_{h}I_{h}(x)\,\widehat{a}_{h}(x)}{\sum_{h}I_{h}(x)},\quad\widehat{p}^{\prime}(x)=\exp\{\widehat{a}(x)\}.$ (75)
@@ Appendix G Baseline Density Estimators
kNN density [Loftsgaarden and Quesenberry, 1965].
@@
$\widehat{p}_{\text{kNN}}(x)$ $=\;\frac{k}{B\,V_{d}\,\big(r_{k}(x)\big)^{d}}.$ (76)
Gaussian KDE with Silverman’s rule [Silverman, 1986].
@@
$\widehat{p}_{\text{KDE}}(x)$ $=\frac{1}{B}\sum_{j=1}^{B}\frac{\exp\!\Big(-\frac{\|x-x^{(j)}\|_{2}^{2}}{2h^{2}}\Big)}{(2\pi)^{d/2}h^{d}},$ (77)
$h$ $=\sigma\left(\frac{4}{d+2}\right)^{\!\frac{1}{d+4}}\,B^{-\frac{1}{d+4}},$ (78)
$\sigma$ $=\tfrac{1}{d}\sum_{\ell=1}^{d}\text{std}(X_{:\ell}).$ (79)
Multivariate Gaussian fit (MGF) [Bishop and Nasrabadi, 2007].
@@
$\quad\log\widehat{p}_{\text{MGF}}(x^{(i)})$ (81)
@@ Appendix H Metrics for Ranking
Spearman’s $\rho$ [Spearman, 1987].
Spearman’s $\rho$ assesses the strength of a monotonic relationship between two rankings. Assign 1-based ranks (1 = best) with average-tied ranks and compute the Pearson correlation between predicted and ground-truth ranks:
@@
Graded Average Precision (gAP) [Moffat et al., 2022].
Graded AP extends Average Precision to graded (multi-level) relevance. First, sort the data by predicted scores. Normalize per query to $[0,1]$: shift if any grade is negative; then, if the maximum exceeds $1$, divide all grades by that maximum. Define
$\mathrm{Prec}@i$ $=\frac{\sum_{j=1}^{i}g_{j}}{i},\quad\mathrm{gAP}=\frac{\sum_{i=1}^{n}g_{i}\,\mathrm{Prec}@i}{\sum_{i=1}^{n}g_{i}}.$ (83)
When $g_{i}\in\{0,1\}$, $\mathrm{gAP}$ reduces to standard AP. Larger values indicate that higher-graded items are concentrated near the top of the ranking.
@@ Appendix I Illustrative Example of Joint ODE Trajectories
Figure 5 visualizes joint ODE trajectories on a 2D Gaussian mixture with three components of mixture weights $1/6$, $1/3$, and $1/2$. Under the IID baseline (no diversity objective), trajectories collapse to the highest-weight component, as expected. Introducing a DPP diversity objective yields Non-IID trajectories that collectively cover all three modes, improving coverage under a fixed trajectory budget. Augmenting DPP with our score-based diversity-velocity regularizer (SR) draws trajectories closer to the mode centers while preserving coverage (see the white rectangle). Finally, importance weighting (IW) assigns a weight to each trajectory. In the full method (DPP + SR + IW), the sample closest to the highest-weight component receives the largest weight (white circle) and that closest to the lowest-weight component receives the smallest, correcting the Non-IID sampling bias and yielding an unbiased estimator with respect to the IID target.
Figure 5: Joint ODE trajectories on a three-component Gaussian mixture. All methods start from identical initial states. “IID” is the baseline without a diversity objective and collapses to the highest-weight mode. “Non-IID (DPP)” uses a determinantal point process objective to encourage sample diversity, enabling the three trajectories to discover all three modes. “SR” denotes our score-based diversity-velocity regularization (the soft version is used in this example). Combining DPP + SR pulls samples toward the underlying modes while preserving coverage (white rectangle). “IW” denotes importance-weight estimation. The full method (DPP + SR + IW) further assigns an importance weight to each trajectory; larger markers indicate higher estimated weight (white circles). For visualization, arrow lengths are rescaled nonlinearly while preserving their relative ordering. The background shows the relative value of $\log p(x)$, where $p$ is the target density corresponding to the IID ODE; yellow indicates higher density and purple indicates lower density.